On a clock, the measure of the angle between the hour and minute hands at 5: 15 a.m. is :
A 60 degrees B 67.5 degrees C 75 degrees D 82.5 degrees E none of these

Answer by dabanfield(803)

On a clock, the measure of the angle between the hour and minute hands at 5: 15 a.m. is :
The hour hand is 1/4 the way between 5 and 6. The minute hand is on 3. The angle between the 3 and the 5 is (2/12)*360 = 60. One fourth the distance between 5 and 6 is (1/4)*(1/12)*360 = 7.5.
The angle between the hands then is 60 + 7.5 = 67.5 (B)